OpenStudy (anonymous):

Will give medal! Find the value of x. A. 30 B. 33.3 C. 75 D. 83.3

OpenStudy (anonymous):



Directrix (directrix):

Is the segment with measure x given to be parallel to the segment with length 50?

OpenStudy (anonymous):

I'm not really sure...this is all the information they give me. :/

Directrix (directrix):

60/100 = x/50 3/5 = x /50 x = 30

OpenStudy (anonymous):

Okay, I got it now. Thanks! :)

Directrix (directrix):

You are welcome.